OpenStudy (anonymous):

Find the derivative of f(x) = 2/x at x = -2

OpenStudy (anonymous):

Would it just be x=-1?

OpenStudy (anonymous):

f(x) = 2*x^(-1) now use y=x^n y'=nx^(n-1)

OpenStudy (anonymous):

so -2x^-2?

OpenStudy (anonymous):

yes, now plug x =-2.

OpenStudy (anonymous):

which is -2/x^2 so -1/2?

OpenStudy (anonymous):

correct

OpenStudy (jdoe0001):

right

OpenStudy (anonymous):

Wooo thanks.

OpenStudy (anonymous):

I got another one like it just ant you to check it. do you want me to start a new question or does doing it on here work?

OpenStudy (anonymous):

you can do it here

OpenStudy (anonymous):

-9/x at x=-4 so it would be 9x^-2 so 9/x^2 so 9/16?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry one last questions. for this one

OpenStudy (anonymous):

Find the derivative of f(x) = 7x + 9 at x = 6 When i work it out i get 7x^0 or just 7+9 which is 16 but that isn't an answer choice?

OpenStudy (anonymous):

derivative of a constant is zero

OpenStudy (anonymous):

What do you mean "of a constant?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

no x in there hence it is constant with respect to x (the variable which you take the derivative with respect to)

OpenStudy (anonymous):

okay so whenever there is no x it automatically becomes a 0?

OpenStudy (anonymous):

making it 7+0 therefore 7?

OpenStudy (anonymous):

yes (if you take the derivative with respect to x)

OpenStudy (anonymous):

see, derivative with respect to x tells you how the function changes when you change x so if there is no x - no change

OpenStudy (anonymous):

Alright :D thanks a lot man :D i think i'm good for now but if i have any more questions i'll let you know if you dont mind that is :P

OpenStudy (anonymous):

yw:)